Shift report 20020117 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 17 2002 - 00:04:03 EST

  • Next message: Apache: "Shift report 20020117 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Jan 17 00:04:03 2002, Shift: 16:00-24:00
    Supervisor: mads, with: 
    Log book#12 pages 70 - 71
    
    Shift Summary:
    ------------------------------------------------------------
    Because of a magnetproblem on the RHIC, there was no beam on  this shift before 2345, when the beam finaly got cogged.
    
    Taking datarun 6645 now.
    
    Earlier today there was a limitid acces, wich Dana used to move the MRS for slide plate. 
    
    Pedestal runs 6622, 6623, 6624 were taken. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Jan 17 2002 - 00:04:37 EST